# KARMA: Efficient Structural Defect Segmentation via Kolmogorov-Arnold Representation Learning

### v1

@@ Abstract
Semantic segmentation of structural defects in civil infrastructure remains challenging due to variable defect appearances, harsh imaging conditions, and significant class imbalance. Current deep learning methods, despite their effectiveness, typically require millions of parameters, rendering them impractical for real-time inspection systems. We introduce KARMA (Kolmogorov-Arnold Representation Mapping Architecture), a highly efficient semantic segmentation framework that models complex defect patterns through compositions of one-dimensional functions rather than conventional convolutions. KARMA features three technical innovations: (1) a parameter-efficient Tiny Kolmogorov-Arnold Network (TiKAN) module leveraging low-rank factorization for KAN-based feature transformation; (2) an optimized feature pyramid structure with separable convolutions for multi-scale defect analysis; and (3) a static-dynamic prototype mechanism that enhances feature representation for imbalanced classes. Extensive experiments on benchmark infrastructure inspection datasets demonstrate that KARMA achieves competitive or superior mean IoU performance compared to state-of-the-art approaches, while using significantly fewer parameters (0.959M vs. 31.04M, a 97% reduction). Operating at 0.264 GFLOPS, KARMA maintains inference speeds suitable for real-time deployment, enabling practical automated infrastructure inspection systems without compromising accuracy. The source code can be accessed at the following URL: https://github.com/faeyelab/karma.
Index Terms:
Structural defect segmentation, Kolmogorov-Arnold networks, semantic segmentation, infrastructure inspection, parameter efficiency, real-time deployment.
@@ I Introduction
Automated detection of structural defects in civil infrastructure is essential for public safety and efficient maintenance [1, 2]. The aging global infrastructure makes accurate segmentation of defects such as cracks, fractures, and deformations increasingly critical [3, 4]. Traditional methods relying on manual video inspection are labor intensive, slow, and error-prone [5]. In municipal sewer and culvert systems, missed defects can lead to catastrophic failures, compromising urban safety and water management. Although semantic segmentation methods offer pixel-level defect localization, existing approaches often struggle to balance accuracy with computational efficiency [3, 5].
The accurate segmentation of structural defects presents several challenges [6, 7]. Defects vary widely in size, shape, and appearance, occurring under various environmental conditions [8, 9]. Inspection imagery often suffers from poor lighting, debris obstructions, and inconsistent camera viewpoints [5]. Furthermore, there is a class imbalance, as rare defects such as holes or collapses occur far less frequently than common problems such as surface cracks [10]. Thus, robust and computationally efficient semantic segmentation methods are required for reliable performance in practical scenarios [6, 10].
Deep learning approaches, particularly Fully Convolutional Networks (FCNs) [11] and U-Net-based models [12, 13], have significantly improved segmentation performance through pixel-level predictions and innovations such as skip connections and feature pyramid modules. However, these models frequently involve millions of parameters, which limits deployment in resource-constrained settings such as inspection robots or real-time monitoring systems [14]. Therefore, there is a clear need for segmentation architectures that achieve high accuracy while remaining computationally efficient.
Kolmogorov–Arnold representation learning offers a promising solution to these challenges [15, 16, 17]. According to the Kolmogorov–Arnold theorem, any continuous multivariate function can be represented by compositions of one-dimensional functions, enabling compact and efficient neural network designs known as Kolmogorov-Arnold Networks (KANs) [18, 19, 20]. Recent applications of KAN in medical image segmentation achieved state-of-the-art results [21], highlighting their potential for structural defect segmentation.
@@
  • An efficient inference pipeline suitable for practical deployment, accompanied by detailed analysis of speed-accuracy trade-offs and memory requirements for edge devices.
@@ III Problem Formulation
Let $\mathcal{I}=\{I_{i}\}_{i=1}^{N}$ be a set of $N$ input images, where each image $I_{i}\in\mathbb{R}^{H\times W\times C}$ has height $H$, width $W$, and $C$ channels. For each image $I_{i}$, there exists a corresponding ground-truth segmentation mask $M_{i}\in\{0,1\}^{H\times W\times K}$, where $K$ denotes the number of defect categories including the background class. The objective is to learn a mapping function $f_{\theta}:\mathcal{I}\rightarrow\mathcal{M}$ parameterized by $\theta$ that can effectively segment structural defects.
@@ IV Proposed Kolmogorov-Arnold Representational Mapping Architecture (KARMA)
We propose KARMA, a novel defect segmentation architecture that enhances the AFPN backbone with Kolmogorov-Arnold representation learning through lightweight TiKAN modules as shown in Fig. 1. Unlike U-KAN which utilizes the original KAN layers in UNet, our method incorporates an efficient TiKAN architecture into an AFPN framework. The key innovation lies in our low-rank formulation of the KA representation that significantly reduces parameters while maintaining expressiveness. KARMA consists of three main components: (1) An AFPN backbone for hierarchical feature extraction; (2) Low-rank TiKAN modules that transform features at each scale using efficient KA representation learning; and (3) A decoder that fuses the enhanced multi-scale features for final segmentation.
@@ IV-A2 Spline-Based Nonlinear Transformation
The spline transformation provides additional flexibility in modeling nonlinear relationships. The grid-based approach allows for precise control over the function approximation, while the noise regularization during training helps prevent overfitting and improves generalization. The factorized spline weight matrix maintains parameter efficiency while allowing for rich representational capacity.
@@ IV-B Integration of TiKAN into AFPN
Our KARMA architecture strategically integrates TiKAN modules into the AFPN backbone through a hierarchical feature pyramid design. The architecture consists of bottom-up and top-down pathways with TiKAN enhancement at critical stages. The first component addresses the challenge of efficient feature extraction through InceptionSepConv blocks, while the second component leverages TiKAN for enhanced representation learning. The multi-scale design allows the network to capture both fine-grained defect details and global contextual information.
@@ IV-B3 Top-down Pathway with Feature Fusion
The top-down pathway implements a progressive feature fusion strategy through depthwise-separable convolutions and upsampling operations, expressed as $\mathbf{p}_{i}=\mathcal{D}(\mathbf{c}_{i})+\mathcal{U}_{2}(\mathbf{p}_{i+1})$ for $i\in\{4,3,2\}$, where $\mathcal{D}$ represents a depthwise-separable convolution with kernel size $1$, and $\mathcal{U}_{2}$ denotes a $2\times$ upsampling operation. This design choice significantly reduces computational cost while maintaining effective feature transformation capability. The additive connections allow the network to combine high-level semantic information from deeper layers with lower-level spatial details from shallower layers, creating a rich multi-scale representation that preserves both fine-grained features and broader contextual information.
@@ IV-B5 Static Dynamic Prototype Mechanism
The static part of this mechanism involves creating representative ‘prototypes’ for each defect class. These prototypes are learned feature vectors that represent the key characteristics and variations within each category. During training, the network adjusts its learned feature representations to align with these class-specific prototypes. This process helps stabilize the feature space, providing reference points that help the model form distinct clusters for each class, even for those with few training examples.
The dynamic part describes how these prototypes are used during network operation, in both training and inference. Instead of only feed-forward classification, the dynamic mechanism allows an adaptive interaction between incoming feature maps and the learned prototypes. This interaction can involve the network adjusting its feature processing based on similarity to the prototypes. For example, features from unclear regions can be made more distinct by relating them to their most relevant prototype, which increases their discriminative power. This dynamic adjustment is important for improving segmentation precision, especially at boundaries, and for consistent performance across all defect types, including rare or subtle ones.
@@ IV-C3 Optimization Strategy
Our optimization strategy involves a single-stage training for 50 epochs using the AdamW optimizer (weight decay of 1e-5, initial learning rate 0.001) combined with a cosine annealing scheduler reducing the learning rate to 1e-6. Data is processed in batches of 16 with gradient clipping (maximum norm 1.0) for training stability. The loss function integrates weighted cross-entropy loss, Dice loss, and Focal loss (weights 0.5, 0.3, and 0.2 respectively), effectively managing class imbalance and ensuring precise boundary representation. The best-performing model is selected based on validation IoU scores to prevent overfitting. Training is conducted in PyTorch on an NVIDIA A100 GPU, ensuring stable convergence and high segmentation accuracy.
@@ V-A1 Datasets
We assess our proposed method using two challenging structural defect datasets. The Structural Defects Dataset (S2DS) [37] contains 743 high-resolution images of concrete surfaces collected via DSLR cameras, mobile phones, and drones. These images are pixel-wise annotated across seven distinct classes: background, cracks (linear fractures), spalling (surface detachment), corrosion (rust), efflorescence (chemical deposits), vegetation (plant growth), and control points (fiducial markers). The dataset is divided into training (563 images), validation (87 images), and testing (93 images) subsets. S2DS is characterized by significant class imbalance (dominated by background pixels), diverse imaging conditions, and subtle textural variations among defects, providing a rigorous benchmark for evaluating model robustness and generalization. Additionally, the Culvert-Sewer Defects Dataset comprises around 6,300 frames [34] extracted from 580 annotated underground inspection videos supplied by the U.S. Army Corps of Engineers and an industry partner. This dataset is split into 70% training, 15% validation, and 15% testing subsets. Frames are manually annotated according to NASSCO PACP standards, covering eight defect classes: cracks, roots, holes, joint issues, deformation, fracture, encrustation/deposits, and loose gasket. These annotations represent real-world scenarios with substantial class imbalance, variable materials, challenging imaging conditions, and multi-scale defect features, making it ideal for testing KARMA’s effectiveness in realistic infrastructure inspection contexts.
@@ V-B KARMA Model Variants
To provide a comprehensive evaluation across different computational budgets and performance targets, we introduce three variants of our KARMA architecture: the baseline KARMA, KARMA High, and KARMA Flash. These variants differ primarily in their channel dimensions throughout the network, the complexity of the KAN module application, and the channel widths in the Feature Pyramid Network (FPN) stages. The motivation behind these variants is to explore the trade-offs between model capacity, computational efficiency (parameters, FLOPs, inference speed), and segmentation accuracy.
@@
KARMA High: This variant is designed to maximize segmentation performance, potentially at the cost of increased computational resources, with the aim of establishing an upper bound on the performance achievable with the KARMA architectural principles by scaling up its capacity. The main differences from the baseline KARMA include increased channel dimensions in the bottom-up InceptionSepConv blocks, resulting in the $c_{5}$ feature map having 1024 channels. Additionally, the KANBlock in the TiKAN Enhancement Module processes these 1024 channels, and the FPN layers in the top-down pathway are wider, operating with 128 channels and utilizing standard convolutions instead of depthwise-separable convolutions for potentially richer feature fusion.
KARMA Flash: This variant is designed for maximum computational efficiency and inference speed, ideal for environments with limited resources or applications requiring real-time processing. It aims to create a lightweight model while preserving the key benefits of the KARMA architecture through several modifications compared to the baseline. These modifications include reducing channel dimensions in the later stages of the bottom-up pathway, with the $c_{5}$ feature map reduced to 384 channels. A pre-KAN $1\times 1$ convolution further decreases the $c_{5}$ features from 384 to 256 channels before they are processed by the KANBlock. The KANBlock in the TiKAN Enhancement Module operates on these reduced 256 channels. Within the KANLayer, the hidden feature dimension is halved relative to its input feature dimension, further reducing parameters in the KAN operations, and the FPN layers in the top-down pathway are narrower, operating with only 32 channels.
@@ V-F Efficiency-Performance Analysis
1. Low-rank TiKAN modules: The TiKAN modules with low-rank factorization efficiently capture complex defect features while reducing parameter counts by over 90% compared to conventional KAN implementations. For instance, as shown in Tables III and IV, removing Low-Rank Adaptation (KARMA (w/o LRA)) significantly increases parameters (e.g., from 0.959M to 5.574M on CSDD) while maintaining similar performance, highlighting the crucial role of LRA in parameter reduction. By representing features within a low-rank subspace, KARMA maintains essential expressiveness required for precise segmentation, eliminating unnecessary parameters.
@@ VI Conclusion
This paper introduced KARMA, an efficient semantic segmentation framework tailored for structural defect detection in infrastructure inspection. KARMA integrates Kolmogorov-Arnold representation learning into an enhanced Feature Pyramid Network, achieving competitive segmentation accuracy with significantly fewer parameters (0.959M) and low computational complexity (0.264 GFLOPS), reducing complexity by approximately 97% compared to conventional models. Evaluations on two challenging defect datasets demonstrated KARMA’s effectiveness, showing strong segmentation performance alongside substantial computational efficiency. Ablation studies confirmed the key contributions of the low-rank TiKAN modules and optimized feature extraction approach to these results. Future research could further optimize TiKAN modules through adaptive rank selection, incorporate selective attention mechanisms, explore applicability to robot-based real-time deployment scenarios, and extend KARMA’s approach to other image analysis tasks.

### v2

@@ Abstract
Semantic segmentation of structural defects in civil infrastructure remains challenging due to variable defect appearances, harsh imaging conditions, and significant class imbalance. Current deep learning methods, despite their effectiveness, typically require millions of parameters, rendering them impractical for real-time inspection systems. We introduce KARMA (Kolmogorov-Arnold Representation Mapping Architecture), a highly efficient semantic segmentation framework that models complex defect patterns through compositions of one-dimensional functions rather than conventional convolutions. KARMA features three technical innovations: (1) a parameter-efficient Tiny Kolmogorov-Arnold Network (TiKAN) module leveraging low-rank factorization for KAN-based feature transformation; (2) an optimized feature pyramid structure with separable convolutions for multi-scale defect analysis; and (3) a static-dynamic prototype mechanism that enhances feature representation for imbalanced classes. Extensive experiments on benchmark infrastructure inspection datasets demonstrate that KARMA achieves competitive or superior mean IoU performance compared to state-of-the-art approaches, while using significantly fewer parameters (0.959M vs. 31.04M, a 97% reduction). Operating at 0.264 GFLOPS, KARMA maintains inference speeds suitable for real-time deployment, enabling practical automated infrastructure inspection systems without compromising accuracy. Real-world hardware validation on NVIDIA Jetson AGX Orin confirms KARMA’s practical deployment capabilities, achieving consistent real-time performance in controlled laboratory environments. The source code can be accessed at https://github.com/faeyelab/karma.
@@ I Introduction
Automated detection of structural defects in civil infrastructure is essential for public safety and efficient maintenance [1, 2]. The aging global infrastructure makes accurate segmentation of defects such as cracks, fractures, and deformations increasingly critical [3, 4]. Traditional methods relying on manual video inspection are labor intensive, slow, and error-prone [5]. In municipal sewer and culvert systems, missed defects can lead to catastrophic failures, compromising urban safety and water management. Although semantic segmentation methods offer pixel-level defect localization, existing approaches often struggle to balance accuracy with computational efficiency [5].
The accurate segmentation of structural defects presents several challenges [6, 7]. Defects vary widely in size, shape, and appearance, occurring under various environmental conditions [8, 9]. Inspection imagery often suffers from poor lighting, debris obstructions, and inconsistent camera viewpoints [5]. Furthermore, there is a class imbalance, as rare defects such as holes or collapses occur far less frequently than common problems such as surface cracks [10]. Thus, robust and computationally efficient semantic segmentation methods are required for reliable performance in practical scenarios [11].
Deep learning approaches, particularly Fully Convolutional Networks (FCNs) [12] and U-Net-based models [13], have significantly improved segmentation performance through pixel-level predictions and innovations such as skip connections and feature pyramid modules. However, these models frequently involve millions of parameters, which limits deployment in resource-constrained settings such as inspection robots or real-time monitoring systems [14]. Therefore, there is a clear need for segmentation architectures that achieve high accuracy while remaining computationally efficient.
Kolmogorov–Arnold representation learning offers a promising solution to these challenges [15, 16, 17]. According to the Kolmogorov–Arnold theorem, any continuous multivariate function can be represented by compositions of one-dimensional functions, enabling compact and efficient neural network designs known as Kolmogorov-Arnold Networks (KANs) [18, 19]. Recent applications of KAN in medical image segmentation achieved state-of-the-art results [20], highlighting their potential for structural defect segmentation.
@@
  • An efficient inference pipeline suitable for practical deployment, accompanied by detailed analysis of speed-accuracy trade-offs, memory requirements for edge devices, and comprehensive real-world hardware validation demonstrating deployment readiness on resource-constrained platforms.
@@ III Problem Formulation
Let $\mathcal{I}=\{I_{i}\}_{i=1}^{N}$ be a set of NN input images, where each image $I_{i}\in\mathbb{R}^{H\times W\times C}$ has height $H$, width $W$, and $C$ channels. For each image $I_{i}$, there exists a corresponding ground-truth segmentation mask $M_{i}\in\{0,1\}^{H\times W\times K}$, where $K$ denotes the number of defect categories including the background class. The objective is to learn a mapping function $f_{\theta}:\mathcal{I}\rightarrow\mathcal{M}$ parameterized by $\theta$ that can effectively segment structural defects.
@@ IV Proposed Kolmogorov-Arnold Representational Mapping Architecture (KARMA)
We introduce KARMA, a defect segmentation architecture improving the AFPN backbone with Kolmogorov-Arnold representation via TiKAN modules. Unlike U-KAN’s original KAN layers in UNet, KARMA integrates an efficient TiKAN design within an AFPN framework. Our innovation uses a low-rank KA representation to reduce parameters while preserving expressiveness. KARMA comprises: (1) an AFPN backbone for hierarchical feature extraction; (2) low-rank TiKAN modules for efficient KA representation at each scale; and (3) a decoder that combines enhanced multi-scale features for segmentation.
@@ IV-A2 Spline-Based Nonlinear Transformation
Spline transformation enhances flexibility in modeling nonlinear relationships. A grid-based approach ensures precise function approximation, and noise regularization prevents overfitting and boosts generalization. The factorized spline weight matrix offers parameter efficiency with rich representation.
@@ IV-B Integration of TiKAN into AFPN
The KARMA architecture integrates TiKAN modules into the AFPN backbone using a hierarchical feature pyramid. It features bottom-up and top-down pathways with TiKAN enhancements at key stages. InceptionSepConv blocks ensure efficient feature extraction, while TiKAN enhances representation learning. This multi-scale design captures fine-grained defect details and global context.
@@ IV-B3 Top-down Pathway with Feature Fusion
The top-down pathway implements a progressive feature fusion strategy through depthwise-separable convolutions and upsampling operations, expressed as $\mathbf{p}_{i}=\mathcal{D}(\mathbf{c}_{i})+\mathcal{U}_{2}(\mathbf{p}_{i+1})$ for $i\in\{4,3,2\}$, where $\mathcal{D}$ represents a depthwise-separable convolution with kernel size 11, and $\mathcal{U}_{2}$ denotes a $2\times$ upsampling operation. This design choice significantly reduces computational cost while maintaining effective feature transformation capability. The additive connections allow the network to combine high-level semantic information from deeper layers with lower-level spatial details from shallower layers, creating a rich multi-scale representation that preserves both fine-grained features and broader contextual information.
@@ IV-B5 Static Dynamic Prototype Mechanism
The static mechanism involves creating prototypes for each defect class, consisting of feature vectors that capture essential characteristics of each category. During training, the network aligns its features with these prototypes to stabilize the feature space and form distinct clusters for each class, even with limited examples.
The dynamic part outlines the use of prototypes during network operation for both training and inference. Instead of simple feed-forward classification, the dynamic mechanism adaptively interacts between incoming feature maps and learned prototypes. The network adjusts based on feature-prototype similarity, enhancing unclear feature distinction and increasing discriminative power. This adjustment boosts segmentation precision, especially at boundaries, ensuring consistent performance across all defect types, including rare or subtle ones.
The static-dynamic prototype mechanism enhances learning by dynamically using reference points in the feature space, allowing KARMA to achieve high segmentation accuracy and stability, even with class imbalances and diverse defect appearances in real-world challenges.
@@ IV-C3 Optimization Strategy
Our optimization strategy uses AdamW for single-stage training over 50 epochs with an initial learning rate of 0.001. A cosine scheduler reduces the learning rate to 1e-6. Batch size is 16, and gradient clipping (max norm 1.0) is applied for stability. The loss function combines weighted cross-entropy, Dice, and Focal losses (weights 0.5, 0.3, 0.2) for class balance and boundary accuracy. Model selection is based on validation IoU scores to prevent overfitting. Training is done in PyTorch on an NVIDIA A100 GPU for stable convergence and high accuracy.
@@ V-A1 Datasets
We evaluate our method with two challenging structural defect datasets. The Structural Defects Dataset (S2DS) [37] features 743 high-res images of concrete surfaces from DSLR cameras, mobile phones, and drones, annotated pixel-wise into seven classes: background, cracks, spalling, corrosion, efflorescence, vegetation, and control points. It is divided into 563 training, 87 validation, and 93 testing images, marked by class imbalance, varied imaging conditions, and subtle defect textures, offering a thorough test for model robustness. The Culvert-Sewer Defects Dataset, from 580 annotated inspection videos provided by the U.S. Army Corps of Engineers and an industry partner, consists of 6,300 frames, split into 70% training, 15% validation, and 15% testing [38]. It follows NASSCO PACP standards for eight defect classes: cracks, roots, holes, joint issues, deformation, fracture, encrustation, and loose gasket. These annotations reflect real-world scenarios with class imbalance, varied materials, challenging conditions, and multi-scale defect features, making it ideal for evaluating KARMA’s effectiveness in infrastructure inspection.
@@ V-B KARMA Model Variants
We introduce three KARMA architecture variants—baseline KARMA, KARMA High, and KARMA Flash—to evaluate different computational budgets and performance goals. These variants differ in channel dimensions, KAN module complexity, and FPN channel widths, aiming to balance model capacity, computational efficiency, and segmentation accuracy.
@@
KARMA High: This variant maximizes segmentation performance by increasing computational resources, aiming to establish an upper performance limit with KARMA principles. Key differences from baseline KARMA involve larger channel dimensions in bottom-up InceptionSepConv blocks, creating a $c_{5}$ feature map with 1024 channels. The KANBlock in the TiKAN Enhancement Module also processes these channels, while FPN layers in the top-down pathway use 128 channels and standard convolutions for richer feature fusion.
KARMA Flash: This variant focuses on computational efficiency and speed, suitable for limited-resource or real-time applications. It creates a lightweight model by modifying the KARMA architecture. Channel dimensions in the bottom-up pathway are reduced to 384 channels, and a pre-KAN convolution further decreases them to 256 channels for KANBlock processing. The KANLayer halves the hidden feature dimension to minimize parameters, and the FPN layers in the top-down pathway are narrower, with 32 channels.
@@ V-F Efficiency-Performance Analysis
1. Low-rank TiKAN modules: TiKAN modules with low-rank factorization capture complex defects efficiently and cut parameters by over 90% compared to regular KANs. As Tables III and IV show, removing Low-Rank Adaptation (KARMA without LRA) raises parameters significantly (e.g., from 0.959M to 5.574M on CSDD) while keeping similar performance, proving LRA’s key role in reducing parameters. KARMA sustains necessary expressiveness for precise segmentation by using a low-rank subspace and eliminating excess parameters.
@@ V-H9 Performance Summary
V-I Real-World Hardware Validation
To validate the practical applicability of KARMA, we conducted a series of experiments on a real-world hardware platform. These experiments were designed to simulate the conditions of a real-world inspection scenario and to verify the performance of our model in a resource-constrained environment. Our deployment of KARMA on an NVIDIA Jetson AGX Orin, a common platform for edge computing in robotics, represents a significant step in transitioning our research from a laboratory setting to real-world applications. Our field tests confirmed that the system could operate as designed under real-world conditions, demonstrating that lightweight deep learning models such as KARMA can be effectively used for in-situ infrastructure inspection.
V-I1 Deployment Validation
To comprehensively validate KARMA, we performed both controlled laboratory experiments and real-world field tests. As shown in Figure 5, our laboratory validation setup used high-resolution printed images of real culvert defects, which were affixed to the walls of the laboratory to simulate a real pipe-infrastructure inspection setting. This testing approach allows for reproducible evaluation of the model’s performance in a controlled environment while maintaining the visual complexity of real-world defects.
Figure 5: Experimental validation of KARMA deployment on a real-world hardware platform.
The experimental setup demonstrates the effective integration of KARMA with a real-time monitoring interface, detailing system health, including computational load and memory usage. The ability to monitor the dual streams of real-time camera images and the processed segmentation output simultaneously allows for necessary quality control and oversight during autonomous inspection missions. The hardware platform was capable of performing real-time defect analysis while maintaining consistent navigation and operational status.
Following the successful laboratory tests, full-scale field trials of KARMA were conducted in realistic pipeline conditions. These trials confirmed the system’s capability in dynamic settings and validated its real-time computational efficiency and accuracy. The inspection results generated by KARMA can be used to create automated reports, such as those in the NASSCO PACP format, which include the type, location, and severity of each defect, providing a comprehensive and actionable assessment of the infrastructure’s condition.
@@ VI Conclusion
This paper introduced KARMA, an efficient semantic segmentation framework tailored for structural defect detection in infrastructure inspection. KARMA integrates Kolmogorov-Arnold representation learning into an enhanced Feature Pyramid Network, achieving competitive segmentation accuracy with significantly fewer parameters (0.959M) and low computational complexity (0.264 GFLOPS), reducing complexity by approximately 97% compared to conventional models. Real-world hardware validation on NVIDIA Jetson AGX Orin demonstrates KARMA’s practical deployment capabilities, confirming its suitability for resource-constrained inspection scenarios. Evaluations on two challenging defect datasets demonstrated KARMA’s effectiveness, showing strong segmentation performance alongside substantial computational efficiency. Ablation studies confirmed the key contributions of the low-rank TiKAN modules and optimized feature extraction approach to these results. The comprehensive hardware validation in laboratory experiments establishes KARMA’s readiness for real-world deployment in infrastructure inspection applications. Future research could further optimize TiKAN modules through adaptive rank selection, incorporate selective attention mechanisms, and extend KARMA’s approach to other image analysis tasks.